Which of the following is the equation in slope-intercept form of the line that passes through (0, 2) and has a slope of -3.
Found 2 solutions by uma, rapaljer:
Answer by uma(370)

The equation of the line through (x1,y1) and slope 'm' is given by:
y - y1 = m(x - x1)
Here m = -3 and x1 = 0, y1 = 2
Thus the equation is y - 2 = -3(x - 0)
==> y - 2 = -3x
==> y = -3x + 2 [adding 2 to both the sides]
This is in slope intercept form
Good Luck!!!

Answer by rapaljer(4551)

y=mx+b, where m is the slope and b is the y intercept.

y =-3x+2

R^2 Retired from SCC